Systems Staff Meeting Notes 11/20/00

Systems Staff Meeting Notes
November 20, 2000

Present: Rich Bennett, Cynthia Bowker, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Geoffrey Stewart, Ying Tang
Announcements
  1. Geoffrey's birthday was Saturday.
  2. Ordered Items Checklist (located in the public folders) has been resurrected by Andy. With his schedule so sporadic for the rest of this semester, this will be a way to keep track of what he has ordered. It will show orders that are in process, not started & completed. Suzy will give permissions to those staff that should have the ability to update the Checklist.
Review Minutes of 11/13/00
  1. Liaisons - Wednesday's training session went well. Feedback has been very positive.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. New Dell order - Bobbie is working on updating the CPU list. The current order is for staff, the next one will be for public machines.
  2. Net Agent - the machines that will run Net Agent will require a lot of resources. Supervisors will need newer, enhanced machines to load the software. Anyone who is on Net Agent will need 128M RAM and at least 200 processor speed.
  3. Barcode Readers - nothing done yet about a trade.
  4. The Net Agent server has been ordered, but will not be here in time to set up, install and go through the checklist before the training period, Dec. 5-7. Will got the other server (DLC) up and running on Friday. We can use it for the Net Agent training. Will need the SQL stuff set up on it - Winston has the specs.
Liaisons
  1. Training sessions complete & successful.
Networking
  1. NT8 - upgrade attempted over the weekend, but it was not successful. Maybe over the Thanksgiving weekend?
  2. Proxy server - Debra notes that we average about 2 calls per week. Fix is to reboot.
  3. Public Printing - still waiting for Xerox to return the server.
  4. Wiring - room 200 still waiting on phones. Six drops in DLC have not been pulled. Debra straightened out the phones in DLC - 3 new extensions have been added.
  5. South Tower - Digital pulled lines on Thursday/Friday. No punch down yet, Bobbie to do it today.
  6. Journalism - switched over to the library server on Friday. NERDC tripped the alarm in Journalism on Sunday. Priscilla West had to come in to reset it. Leilani called on Friday about the subnet masks in Journalism ending in 128 - she didn't want to take Geoffrey's recommendation on the settings. Circ IDs will also be a problem. Leilani & Marvin are out this week; Gary Cornwell will be the contact for Journalism.
Software Projects Status
  1. Film Log - Preservation is ready to transfer the data, but Hilton is out until next Tuesday.
  2. FULoader/MARCon mods - Winston has been making small fixes; project still ongoing.
  3. Facilities Access Inventory on NT & Key Fob program - Winston has been talking with Andy Barclay from Semco about how to incorporate their backup with our server.
  4. IRDB - still ongoing.
  5. Org chart - IE security zones giving some problems, Cynthia & Winston are investigating.
  6. Binding database - not assigned.
  7. Mellon updates are complete.
DP Coordinator
  1. Andy will be working twice a week between now and the spring term. He should be back to at least four days a week by next semester.
  2. Debra has a lead on a possible replacement for Jorga - a student that works for Geoffrey's wife is looking for an additional job.
Calendar
  1. Will & Hilton both out Nov. 20-22 & 27.
  2. Martha out Nov. 22, afternoon.
  3. Geoffrey out November 22.
  4. Bobbie out November 27.
  5. Cynthia out Dec. 1, 8 & 15.
  6. Bobbie out Dec. 1, 4, 8, 11, 15, 18, 22 & 26.
  7. Martha out Dec. 7-8.
  8. Debra out December 20-27.
  9. Martha out Dec. 26-Jan. 2.
  10. Bill returns on Jan. 8.


Back to the Systems Homepage

Last updated November 27, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu